"A body is projected upward with velocity 20m/s from the top of a cliff of height 160m. in what times will it reach the ground? (g=10m/s)"

 

 

Grade:undefined

2 Answers

Neeti
8 years ago
for upward journey, initial velocity = 20m/s and final is 0  v= u + at gives you the time for upward journey.
Now calculate the distance travelled by the body during this upward journey. The distance s plus 160 will be the distance covered by the body during downward journey and the initial velocity for the downward journey shall be zero. using s = 1/2at2 calculate time for downward journey and add the two. 
Please try it once and if you’re unable to get the answer i’d be happy to help :)
 
Please approve the answer if you feel it’s useful and if you like it :)
Vaibhav Gupta
8 years ago
u=20m/s
v=0m/s
displacement = -160m
g = 10m/s
 
-160 = 20t – 5t2
5t2 – 20t – 160 = 0
 
solving the above equation we get
t = 8 or t = -4
 
since t cannot be negative
therefore, t = 8 sec
 
PLEASE APPROVE THE ANSWER IF IT HELPS
THANKS